OpenStudy (anonymous):

(3 over 5√15)(16√5) simplify

OpenStudy (ajprincess):

√15=√3*√5 3=√3*√3 can u do nw usin these?

OpenStudy (anonymous):

im horrible with square roots....

OpenStudy (ajprincess):

 is it clear?